OpenStudy (anonymous):

Solve the system of equations using substitution. y = 2x + 3 y = 3x + 1

hartnn (hartnn):

so you can substitute y from first equation in 2nd equation, 2x+3 = 3x+1 can you get x from here ?

OpenStudy (anonymous):

no

hartnn (hartnn):

what problem are you facing ? start by subtracting 2x from both sides.

OpenStudy (anonymous):

2x+3=3x+1 x=2 by put in given equcation y=7